# Differential Test Performance and Peer Effects

### v1

@@ Appendix B Auxiliary Lemmas
By Assumptions 3, 4 and 6, $\left(n_{c}-1-\rho\right)/\left(n_{c}-1\right)$, $\left(1+\rho\right)$, $\gamma_{j}^{2}$, $\rho_{0,\tau_{c}}^{2}$, and $\sigma_{0,t}^{2}$ in (39), (40), (41) are positive and uniformly bounded away from 0. Hence by Lemma B.1(v), $I+\rho M$, $\Omega(\gamma)$ and $\Sigma_{t}$ are non-singular. An inspection of Equations (38)-(41) shows that all matrices in the set can be written in the form of $\operatorname{diag}_{c=1}^{C}\left(p_{c}I_{c}^{\ast}+q_{c}J_{c}^{\ast}\right),$ with $p_{c}$ and $q_{c}$ continuously differentiable in $\theta$ and $\gamma$ and uniformly bounded. Also note that by Assumption 5, the row and column sums of $A$ are uniformly bounded in absolute value. Utilizing remark A.1 in Kelejian and Prucha (2004) the corollary thus follows from Lemma B.1. ∎
We use these properties frequently in the derivations that follow.
Lemma B.2.
Suppose Assumptions 3, 4, 5, and 6 hold. Consider the block diagonal matrix $\mathcal{M}\left(\rho\right)=\operatorname{diag}{}_{c=1}^{C}\left(\mathcal{M}_{c}\left(\rho\right)\right)$ with $\mathcal{M}_{c}\left(\rho\right)=\left(I_{c}+\rho M_{c}\right)^{-1}\left(I_{c}+\rho_{0}M_{c}\right)$. Then for all $\gamma\in\Gamma$,
and
Proof.
Noting that $\operatorname{tr}\left(A_{c}\right)=0$ and hence $\operatorname{tr}\left(A_{c}I_{c}^{\ast}\right)=\operatorname{tr}\left(A_{c}(I-J_{c}^{\ast})\right)=-\operatorname{tr}\left(A_{c}J_{c}^{\ast}\right)$, and in light of Lemma B.1
We have
where
Hence
Next observe that $\operatorname{tr}\left(A_{c}J_{c}^{*}\right)=\mathbf{1}_{c}^{{}^{\prime}}A_{c}\mathbf{1}_{c}/n_{c}\geqslant\text{$\underline{K}$}_{a}>0$ and that for any $n_{c}\geqslant 2$ and $-1<\rho<1$, both $n_{c}-1-\rho$ and $1+\rho$ are positive, and therefore $C(\rho,n_{c})$ is positive for all $\rho\in\left(-1,1\right)$. Therefore, $\operatorname{tr}\left[\mathcal{M}_{c}\left(\rho\right)A_{c}\mathcal{M}_{c}\left(\rho\right)\right]<0$ if $\rho>\rho_{0}$, $\operatorname{tr}\left[\mathcal{M}_{c}\left(\rho\right)A_{c}\mathcal{M}_{c}\left(\rho\right)\right]>0$ if $\rho<\rho_{0}$ and $\operatorname{tr}\left[\mathcal{M}_{c}\left(\rho\right)A_{c}\mathcal{M}_{c}\left(\rho\right)\right]$ has the same sign for all $c$. In all, for any value of $\gamma$ the only solution for
is $\rho=\rho_{0}$. Furthermore, by Assumptions 4 and 6 there exists some constants $c_{D}$ and $c_{\gamma}$ such that for any $\rho_{0}$ in the interior of the parameter space, $C(\rho,n_{c})>c_{D}>0$ and $\frac{\gamma_{\tau_{c},0}^{2}}{\gamma_{\tau_{c}}^{2}}>c_{\gamma}>0$. This implies that,
$\left|\operatorname{tr}(D(\rho))\right|\geqslant|\rho_{0}-\rho|\sum_{c=1}^{C}\text{$\underline{K}$}_{a}c_{D}c_{\gamma},$
$\left|\frac{1}{n}\operatorname{tr}(D(\rho))\right|\geqslant|\rho_{0}-\rho|\text{$\underline{K}$}_{a}c_{D}c_{\gamma}\frac{C}{n}.$
Lemma B.3.
Let Assumption 2 hold. Then, $u_{it}$ defined in Theorem 2.1 is a martingale difference sequence w.r.t to the filtration $\mathcal{A}_{n,\nu}$ with $\mathcal{A}_{n,0}=\mathcal{Z}_{n}^{*}=\sigma\left(S,\alpha,X,z,\zeta^{*}\right)$ and
$\mathcal{A}_{n,\left(t-1\right)n+i-1}=\sigma\left(S,\alpha,X,z,\zeta^{*},\left\{u_{j,t-1}\right\}_{j=1}^{n},\left\{u_{jt}\right\}_{j=1}^{i-1}\right)\text{\,for $t=1,2;i=1,\ldots,n$}$ (43)
Proof.
Appendix C Proofs
Proof of Theorem 2.1.
By the assumption on the boundedness of the ($4+\eta$)-th moments all considered expectations and conditional expectations exist. Since $u_{it}=y_{it}^{*}-E\left[y_{it}^{*}|\mathcal{F}_{n,i,t}\right]$ we have $E\left[u_{it}|\mathcal{F}_{n,i,t}\right]=0$ by construction. Observing that $E\left[y_{it}^{*}|\mathcal{F}_{n,i,t}\right]=\kappa_{i}f_{t}$ and that $\kappa_{i}$ is $\mathcal{Z}_{n}^{*}$-measurable it follows that conditional on $\mathcal{Z}_{n}^{*}$ knowledge of $y_{it}^{*}$ is equivalent to knowledge of $u_{it}$. To see this let $u_{-i,t}=\left(u_{1t},..,u_{i-1,t},u_{i+1,t},...,u_{n,t}\right)^{\prime}$ and use the notation $u_{-t}$ for the unobserved components related to the other test. Define $U=\left(S,\alpha,X,z,\zeta^{*},u_{-t},u_{-i,t},\kappa_{i},f_{t}\right)$ assuming for simplicity of notations that vectors are row vectors, and let $V=\left(S,\alpha,X,z,\zeta^{*},y_{-t}^{*},y_{-i,t}^{*},\kappa_{i},f_{t}\right)$. Then $V=g(U)$ where $g(.)$ is Borel measurable and thus $\sigma(V)\subseteq\sigma(U)$. Analogously $U=h(V)$ where $h(.)$ is Borel measurable and thus $\sigma(U)\subseteq\sigma(V)$. Consequently $\sigma(U)=\sigma(V)$. Since $\kappa_{i},f_{t}$ are measurable w.r.t. $\sigma(S,\alpha,X,\zeta^{*},z)$ it follows that
$\sigma(S,\alpha,X,z,\zeta^{*},u_{-t},u_{-i,t},\kappa_{i},f_{t})=\sigma(S,\alpha,X,z,\zeta^{*},u_{-t},u_{-i,t})$
and
$\sigma(S,\alpha,X,z,\zeta^{*},y_{-t}^{*},y_{-i,t}^{*},\kappa_{i},f_{t})=\sigma(S,\alpha,X,z,\zeta^{*},y_{-t}^{*},y_{-i,t}^{*}),$
which shows that
$\mathcal{B}_{n,i,t}=\sigma\left(S,\alpha,X,z,\zeta^{*},u_{-t},u_{-i,t}\right)=\sigma\left(S,\alpha,X,z,\zeta^{*},y_{-t}^{*},y_{-i,t}^{*}\right)=\mathcal{F}_{n,i,t}.$
Next observe that the $u_{jt}$ are $\mathcal{F}_{n,i,t}$-measurable for $i\neq j$, and hence by iterated expectations $\operatorname{Cov}\left(u_{it}u_{jt}|\mathcal{F}_{n,i,t}\right)=E\left[u_{jt}E\left(u_{it}|\mathcal{F}_{n,i,t}\right)\right]=0$. Observe further that for $s<t$ the $u_{js}$ are $\mathcal{F}_{n,i,t}$-measurable for all $i$, and hence by iterated expectations $\operatorname{Cov}\left(u_{it}u_{js}|\mathcal{F}_{n,i,t}\right)=E\left[u_{js}E\left[u_{it}|\mathcal{F}_{n,i,t}\right]\right]=0$. The corresponding claims regarding the mean and covariances of the $u_{it}$ conditional on $\mathcal{G}_{n,i,t}\subseteq\mathcal{F}_{n,i,t}$ follow immediately by iterated expectations. This proves the first part of the lemma.
E$\left[y_{ii}^{*}-\kappa_{i}f_{t}|\mathcal{G}_{n,i,t}\right]$ $=$ $E\left[E\left[y_{it}^{*}|\mathcal{F}_{n,i,t}\right]|\mathcal{G}_{n,i,t}\right]-E\left[\kappa_{i}f_{t}|\mathcal{G}_{n,i,t}\right]=E\left[\kappa_{i}f_{t}|\mathcal{G}_{n,i,t}\right]-E\left[\kappa_{i}f_{t}|\mathcal{G}_{n,i,t}\right]=0.$
Thus by the almost sure uniqueness of conditional expectations, Assumption 2 cannot hold if $E\left[y_{it}^{*}-\kappa_{i}f_{t}|\mathcal{G}_{n,i,t}\right]\neq 0$. ∎
Note that Lemma 4.1, Corollary 4.1, and Lemma 4.3 can be viewed as special cases of Lemma 4.2. We therefore prove Lemma 4.2 first.
Proof of Lemma 4.2.
In light of (9) it follows that,
$y_{1}-f_{1}y_{2}-\text{$\underline{X}$}\delta=\mu^{\ast}(f_{1,0}-f_{1})+\underline{X}\left(\delta(f_{1},\rho_{0},\beta_{0})-\delta\right)+(I+\rho_{0}M)(u_{1}-f_{1}u_{2}).$ (46)
Substitution of (46) into (14) yields
$u^{+}(\theta,\gamma)$ $=$ $\Omega(\gamma)^{-1/2}(I+\rho M)^{-1}$
recalling that $V(\rho,\gamma)=\Omega(\gamma)^{-1/2}(I+\rho M)^{-1}$ and $\mathcal{M}(\rho)=(I+\rho M)^{-1}(I+\rho_{0}M)$.
We first analyze the linear moment conditions. Note that
Let
Observe that
where
the r.h.s. is seen to hold in light of (52) and (54), and since $H$ is $\mathcal{Z}_{n}$-measurable and thus
The assumptions imply that $K_{n}$ has full rank a.s. To see this, note that in light of (39), (40), Lemma (B.1)(iv) implies that
$V(\rho,\gamma)=\operatorname{diag}_{c=1}^{C}\left(\frac{1}{\gamma_{\tau_{c}}}\frac{n_{c}-1-\rho}{n_{c}-1}I_{c}^{\ast}+\frac{1}{\gamma_{\tau_{c}}}\left(1+\rho\right)J_{c}^{\ast}\right).$ (58)
Under Assumptions 4 and 6, there exists some constants $c_{e}$ and $C_{e}$ such that $0<c_{e}<\frac{1}{\gamma_{\tau_{c}}}\frac{n_{c}-1-\rho}{n_{c}-1}<C_{e}<\infty$ and $0<c_{e}<\frac{1}{\gamma_{\tau_{c}}}\left(1+\rho\right)<C_{e}<\infty$. In light of Lemma B.1(iii), the eigenvalues of $V(\rho,\gamma)$ are uniformly bounded below by $c_{e}>0$. By Assumption 7 we have $\lambda_{min}\left(n^{-1}\underline{X}^{\prime}\underline{X}\right)\geq\xi_{X}>0$ and thus $\lambda_{min}\left(n^{-1}\underline{X}^{\prime}V(\rho,\gamma)\underline{X}\right)\geq c_{e}\xi_{X}>0$ uniformly for all $\gamma$, $\rho$, and $n$.
Observe furthermore that
$E\left[(V^{1/2}z)^{\prime}Q_{V^{1/2}X}(V^{1/2}y_{2})\left|\mathcal{Z}_{n}\right.\right]$ $=$ $E\left[(V^{1/2}z)^{\prime}Q_{V^{1/2}X}(V^{1/2}\mu^{*})\left|\mathcal{Z}_{n}\right.\right]$
in light of (9) and since $Q_{V^{1/2}X}$ is orthogonal to $V^{1/2}\underline{X}$ and $E\left[u_{i2}|\mathcal{Z}_{n}\right]=0$ by Theorem 2.1. Thus
$\inf_{\gamma\in\Gamma,\rho\in[-K_{\rho},K_{\rho}]}\left|\operatorname{det}\left(E\left[K_{n}(\rho,\gamma)\left|\mathcal{Z}_{n}\right.\right]\right)\right|$$\geq$ $\left(\xi_{X}c_{e}\right)^{p_{x}}\inf_{\gamma\in\Gamma,\rho\in[-K_{\rho},K_{\rho}]}n^{-1}\left|E\left[(V^{1/2}z)^{\prime}Q_{V^{1/2}X}(V^{1/2}y_{2})\left|\mathcal{Z}_{n}\right.\right]\right|$
$\geq$ $\left(\xi_{X}c_{e}\right)^{p_{x}}K_{y}>0$
Once $f_{1,0}$ and $\delta_{0}$ are identified, $\rho$ can be identified from the quadratic moment condition. To see this, note that when evaluated at $\delta_{0}$ and $f_{1,0},$ $u^{+}(\theta,\gamma)$ in Equation (C) becomes
$=\Omega(\gamma)^{-1/2}\mathcal{M}(\rho)(u_{1}-f_{1,0}u_{2}).$
Then,
[…]
Proof of Lemma 4.1 and Corollary 4.1.
Lemma 4.1 is a special case of Lemma 4.2 without covariates. Note that Assumptions 1, 2, 3, 4, 5, and 6 hold. Without covariates Equation (61) reduces to $K_{n}\left(\rho,\gamma\right)=z^{\prime}V\left(\rho,\gamma\right)\mu^{*}$ where $V\left(\rho,\gamma\right)=\Omega\left(\gamma\right)^{-1/2}\left(I+\rho M\right)^{-1}.$
Recall that $z=\left[z_{1}^{\prime},...,z_{c}^{\prime},...,z_{C}^{\prime}\right]^{\prime}$ with $z_{c}=\grave{z}_{c}\mathbf{1}_{c}$. Utilizing Lemma B.1,
leads to
where the inequality follows from the conditions imposed in the lemma, in particular (28). With $\rho\in(-1,1)$, E$\left[K_{n}\left(\rho,\gamma\right)|\mathcal{Z}_{n}\right]\neq 0$ such that (57) only has one solution, $f_{1}=f_{1,0}$. Identification of $\rho_{0}$ from the quadratic moment is not affected by the absence of covariates. Thus Lemma 4.1 follows from Lemma 4.2.
Proof of Lemma 4.3.
Observe that utilizing (45) we have
Observe that $E\left[\tilde{K}_{n}\right]$ has full rank, as is readily seen from the proof of Lemma 4.2 in light of (C) with $V=I$. This implies that the only solution to (64) is $\delta=\delta_{0}$ and $f_{1}=f_{1,0}$. It now follows that
such that the quadratic moment condition is
By Lemma B.2 it follows that the only solution to (65) is $\rho=\rho_{0}.$
Proof of Theorem 4.1.
To prove part (i) of the theorem we establish the conditions of Lemma 3.1 in Pötscher and Prucha (1997).
For the linear moment conditions use (52). We have
@@ Appendix C Proofs
Consider $III$. From Corollary B.1 and Assumption 9, we have $sup_{\theta,\gamma}|III|\rightarrow_{p}0$.
Now consider the quadratic moment conditions. Let $\vartheta=(\delta^{\prime},f_{1})^{\prime}$, $\vartheta_{0}=(\delta^{\prime}(f_{1},\rho_{0},\beta_{0}),f_{1,0})^{\prime}$ and $\Delta_{1}=[\underline{X},\mu^{\ast}]$. We point out that $\vartheta_{0}$ is a function of $f_{1}$. We omit this dependence for notational convenience and because $f_{1}$ is pinned down by linear moment conditions for $f_{1}$ directly, or in other words by equations related to the second component of $\vartheta$. In light of (C) we have
@@
By Corollary B.1 and Assumption 9, $II$ converges uniformly to 0. Consider $III$, let $B(\theta,\gamma)=\mathcal{M}(\rho)^{\prime}\Omega(\gamma)^{-1/2}A\Omega(\gamma)^{-1/2}\mathcal{M}(\rho)$, then
@@
Each term on the R.H.S. converges to its mean uniformly in $\theta$ and $\gamma$ under Corollary B.1 and Assumption 9, and the mean is finite and continuously differentiable in $\theta,\gamma$ Consequently, $\frac{1}{n}(u_{1}-f_{1}u_{2})^{\prime}B(\theta,\gamma)(u_{1}-f_{1}u_{2})$ converges uniformly to the limit of its mean
@@
We proceed to show, building on our results for finite $n$, that $\theta_{0}$ is also the unique solution vector of the limiting moment condition $\mathfrak{m}\left(\theta,\gamma_{*}\right)=0$. Observe that from (54), (66) and (67) we have $\left[\mathscr{U}_{H,x}\left(\theta,\gamma_{*}\right),\mathscr{U}_{H,\mu}\left(\theta,\gamma_{*}\right)\right]=\lim_{n\rightarrow\infty}E\left[K_{n}(\rho,\gamma_{*})\right],$ and by (C) we have $\inf_{\gamma\in\Gamma,\rho}\left|\operatorname{det}\left(E\left[K_{n}(\rho,\gamma)\right]\right)\right|\geq C_{K}>0$ for some $C_{K}>0$ and all $n.$ Since $\operatorname{det}(.)$ is a continuous function it follows that for any $\theta,\gamma_{*}$ we have
@@
By Lemma B.2, $\mathfrak{m}_{q}\left(\rho,f_{1,0},\delta_{0},\gamma_{*}\right)=0$ if and only if $\rho=\rho_{0}$, and consequently $\mathfrak{m}\left(\theta,\gamma_{*}\right)=0$ if and only if $\theta=\theta_{0}$. Assumption 9 implies that $\mathfrak{m}\left(\theta,\gamma_{*}\right)$ and thus $\mathfrak{m}\left(\theta,\gamma_{*}\right)^{\prime}\Xi^{-1}\mathfrak{m}\left(\theta,\gamma_{*}\right)$ is continuous. Since the parameter space is compact it follows that $\theta_{0}=[$$\rho_{0},f_{1,0},\delta_{0}^{\prime}]^{\prime}$ is the identifiable unique minimizer of $\mathfrak{m}\left(\theta,\gamma_{*}\right)^{\prime}\Xi^{-1}\mathfrak{m}\left(\theta,\gamma_{*}\right)$; cp. the discussion after Definition 3.1 in Pötscher and Prucha (1997). Consistency now follows immediately from Lemma 3.1 in Pötscher and Prucha (1997).
@@
[…]
@@
Finally, using Lemma B.2 and Assumption 9 it follows that
@@
Proof of Theorem 4.2.

### v2

@@ Appendix C Consistency of Parameter Estimators
To prove part (i) of the theorem we establish the conditions of Lemma 3.1 in Pötscher and Prucha (1997). For the linear moment conditions use (33). We have
@@
Consider $III$. From Corollary E.1 and Assumption 11, we have $sup_{\theta,\gamma}|III|\rightarrow_{p}0$.
Now consider the quadratic moment conditions. Let $\vartheta=(\delta^{\prime},f_{1})^{\prime}$, $\vartheta_{0}=(\delta^{\prime}(f_{1},\rho_{0},\beta_{0}),f_{1,0})^{\prime}$ and $\Delta_{1}=[\underline{X},\mu^{\ast}]$. We point out that $\vartheta_{0}$ is a function of $f_{1}$. We omit this dependence for notational convenience and because $f_{1}$ is pinned down by linear moment conditions for $f_{1}$ directly, or in other words by equations related to the second component of $\vartheta$. In light of (A) we have
@@
By Corollary E.1 and Assumption 11, $II$ converges uniformly to 0. Consider $III$, let $B(\theta,\gamma)=\mathcal{M}(\rho)^{\prime}\Omega(\gamma)^{-1/2}A\Omega(\gamma)^{-1/2}\mathcal{M}(\rho)$, then
@@
Each term on the R.H.S. converges to its mean uniformly in $\theta$ and $\gamma$ under Corollary E.1 and Assumption 11, and the mean is finite and continuously differentiable in $\theta,\gamma$ Consequently, $\frac{1}{n}(u_{1}-f_{1}u_{2})^{\prime}B(\theta,\gamma)(u_{1}-f_{1}u_{2})$ converges uniformly to the limit of its mean
@@
We proceed to show, building on our results for finite $n$, that $\theta_{0}$ is also the unique solution vector of the limiting moment condition $\mathfrak{m}\left(\theta,\gamma_{*}\right)=0$. Observe that from (35), (52) and (53) we have $\left[\mathscr{U}_{H,x}\left(\theta,\gamma_{*}\right),\mathscr{U}_{H,\mu}\left(\theta,\gamma_{*}\right)\right]=\lim_{n\rightarrow\infty}E\left[K_{n}(\rho,\gamma_{*})\right],$ and by (44) we have $\inf_{\gamma\in\Gamma,\rho}\left|\operatorname{det}\left(E\left[K_{n}(\rho,\gamma)\right]\right)\right|\geq C_{K}>0$ for some $C_{K}>0$ and all $n.$ Since $\operatorname{det}(.)$ is a continuous function it follows that for any $\theta,\gamma_{*}$ we have
@@
By Lemma E.2, $\mathfrak{m}_{q}\left(\rho,f_{1,0},\delta_{0},\gamma_{*}\right)=0$ if and only if $\rho=\rho_{0}$, and consequently $\mathfrak{m}\left(\theta,\gamma_{*}\right)=0$ if and only if $\theta=\theta_{0}$. Assumption 11 implies that $\mathfrak{m}\left(\theta,\gamma_{*}\right)$ and thus $\mathfrak{m}\left(\theta,\gamma_{*}\right)^{\prime}\Xi^{-1}\mathfrak{m}\left(\theta,\gamma_{*}\right)$ is continuous. Since the parameter space is compact it follows that $\theta_{0}=[$$\rho_{0},f_{1,0},\delta_{0}^{\prime}]^{\prime}$ is the identifiable unique minimizer of $\mathfrak{m}\left(\theta,\gamma_{*}\right)^{\prime}\Xi^{-1}\mathfrak{m}\left(\theta,\gamma_{*}\right)$; cp. the discussion after Definition 3.1 in Pötscher and Prucha (1997). Consistency now follows immediately from Lemma 3.1 in Pötscher and Prucha (1997).
@@
where we used (12) for $y_{1}-f_{1,0}y_{2}$. By Assumptions 8 and 11 as well as (13) it follows that $n^{-1}\underline{X}^{\prime}\underline{X}\rightarrow_{p}\mathscr{U}_{XX}$ where $\mathscr{U}_{XX}$ is positive definite and symmetric and $n^{-1}\underline{X}^{\prime}y_{2}\rightarrow_{p}\mathscr{U}_{Xy}$ where $\mathscr{U}_{Xy}$ is bounded. This implies that $II\rightarrow_{p}0.$ For $I$ recall from above that $n^{-1}\underline{X}^{\prime}\left(I_{n}+\rho_{0}M\right)\left(u_{1}-u_{2}f_{1,0}\right)\rightarrow_{p}0.$ This shows that $\tilde{\delta}\rightarrow_{p}\delta_{0}.$
@@
Finally, using Lemma E.2 and Assumption 11 it follows that
@@
Appendix D Asymptotic Normality of Parameter Estimators
Assumption 12.
The first step inefficient but consistent estimators $\tilde{f}_{1}$, $\tilde{\delta}$ and $\tilde{\rho}$ defined in (14), (15) and (16) can be used to obtain consistent estimates of the variance parameters $\gamma.$ Form the residuals
Efficient GMM estimators can now be formed by plugging $\hat{\gamma}$ into $u^{+}(\theta,\gamma)$ defined in (18). Recall that $\Omega(\gamma)=\operatorname{diag}_{c=1}^{C}(\gamma_{\tau_{c}}^{2}I_{c}).$ In the following we denote with $K_{u}$ a generic finite constant (which is taken, w.o.l.o.g., to be greater than one) and which is invariant over $t=1,2$, $i=1,\ldots,n$, $n\geq 1$. The central limit theorem in Proposition 3 of Kuersteiner and Prucha (2020) can be applied to obtain distributional approximations for estimators based on $u^{+}$. The central limit theorem of Kuersteiner and Prucha (2020) requires that $E\left[\left|u_{it}u_{is}\right|^{1+\eta_{u}}|\mathcal{Z}_{n}\right]\leqslant K_{u}$ with $\eta_{u}>0$ for $t,s=1,2\text{ }$ and all $i$. This condition is implied by Assumption 2, as Lemma B.1 demonstrates. When setting
$\tilde{\gamma}=\left(1,1,...,1\right)^{\prime}$ (71)
it follows that $\Omega(\tilde{\gamma})=I_{n}$. The feasible efficient GMM estimator for $\theta_{0}=\left(\rho_{0},f_{1,0},\delta_{0}^{\prime}\right)^{\prime}$ is defined as $\hat{\theta}\left(\hat{\gamma}\right)=\operatorname{argmin}_{\theta}Q_{n}\left(\theta,\hat{\gamma}\right)$ where $Q_{n}\left(\theta,\gamma\right)$ is given in (23) and the inefficient counterpart $\hat{\theta}\left(\tilde{\gamma}\right)$ is defined in an analogous way. The next theorem establishes the limiting distribution of $\hat{\theta}\left(\tilde{\gamma}\right)$ and $\hat{\theta}\left(\hat{\gamma}\right).$ The limiting distribution of $\hat{\theta}\left(\tilde{\gamma}\right)$ for $\tilde{\gamma}$ defined in (71) provides a distributional result for an inefficient GMM estimator. The limiting distribution of the first step estimators defined in (14), (15) and (16) can be deduced from the theorem by working out the limiting distribution of $m_{\epsilon}\left(\theta\right).$ We omit the details.
Theorem D.1.
Let Assumptions 3-12 hold. Let $\hat{\gamma}$ be as defined in (70) and $\tilde{\gamma}$ as defined in (71). For $\gamma_{n}\rightarrow_{p}\gamma$ let $\operatorname{plim}_{n\rightarrow\infty}n^{-1/2}\partial m_{n}\left(\hat{\theta}\left(\gamma_{n}\right),\gamma_{n}\right)/\partial\theta=G(\gamma)$ and $\Xi=\operatorname{plim}_{n\rightarrow\infty}\Xi_{n}$, where $\Xi_{n}$ is defined in (22). Then $\hat{\gamma}$$\rightarrow_{p}\gamma_{0}$, and for $\gamma_{n}=\hat{\gamma}$ or $\gamma_{n}=\tilde{\gamma}$,
In Appendix F we provide explicit formulas for the derivatives $G$. Together with the expressions for $\Xi_{n}$ it is then easy to construct sample based estimators $\hat{G}$ and $\hat{\Xi}$. Letting $\hat{\Psi}_{\gamma}=\left(\hat{G}^{\prime}\hat{\Xi}^{-1}\hat{G}\right)^{-1}$ it follows from a standard Slutsky argument that $\hat{\Psi}_{\gamma}^{-1/2}\sqrt{n}\left(\hat{\theta}\left(\hat{\gamma}\right)-\theta_{0}\right)\rightarrow_{d}N\left(0,I\right).$
Proof of Theorem D.1.